# http://www.michaeltimothybennett.com/

### v3

@@ 3 Limits of Intelligence
As stated earlier, intelligence is often equated with “the ability to generalise”, which amounts to sample efficiency in learning the policy which “caused” examples, and thus predictive accuracy. We wish to understand the objective upper limits of intelligence. It has been formally proven that $<_{w}$ is the optimal choice of proxy for sample efficiency [8, prop. 1, 2, 3] (and weak policies generalised at $110-500\%$ the rate of simple policies in experiments). In other words, the upper bound of intelligent behaviour in the context of an abstraction layer and task is attained by the weakness proxy. To obtain an upper bound on superintelligence, unconstrained by any particular abstraction layer, we need to first define a task “independent” of any particular embodiment. First, we need to formalise a version of the same task across different vocabularies. We use the set of all tasks in all vocabularies $\Gamma=\underset{\mathfrak{v}\subseteq P}{\bigcup}\Gamma_{\mathfrak{v}}$, and the powerset $2^{P}$ of $P$.
Definition 6 ($\lambda$-tasks)
An uninstantiated task is a function $\lambda:2^{P}\rightarrow\Gamma$ that takes a vocabulary $v\in 2^{P}$ and returns a particular task in that vocabulary, such that if $v\subset w$, $\lambda(v)=\alpha$ and $\lambda(w)=\omega$, then $\Pi_{\alpha}\subset\Pi_{\omega}$. An uninstantiated task is a $\mathfrak{v}$-task that is yet to be instantiated in a particular vocabulary.
Next, we need to define how we measure the “utility” of intelligence, or the the extent to which it is possible and advantageous to be sample efficient given a particular abstraction layer or embodiment. This is because more intelligence is not always useful. Intuitively, making a human more intelligent is unlikely to improve their driving. Likewise, intelligence conveys no advantage if one’s task is to interpret random noise.
@@
Every task $\gamma\in\Gamma$ has a “utility” computed as $\epsilon:\Gamma\rightarrow\mathbb{N}$ such that $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert)$.
Definition 8 (upper bound of intelligent behaviour)
Given an uninstantiated task $\lambda$, sample efficiency is maximised if weakness is the proxy and the vocabulary maximises utility. This means choosing a policy $\pi$ in a vocabulary $\mathfrak{v}$ such that if $\lambda(\mathfrak{v})=\alpha$ then there does not exists $w\in 2^{P}$ s.t. $\epsilon(\lambda(w))>\epsilon(\lambda(\mathfrak{v}))$, and $\pi\in\Pi_{\alpha}$, and $<_{w}$ is maximised by $\pi$.
3.0.1 Concluding remarks:
Here we proposed computational dualism, and argued that conceptions of artificial intelligence should be updated to reflect more recent philosophy of mind. We have discussed and expanded upon a formalism of enactivism, which allows for objective claims regarding intelligent behaviour. Variations of these definitions have already been used to establish optimally sample efficiency and causal learning [8, 21]. We have formalised the objective upper bound of intelligent behaviour, but to attain this upper bound one would need to search an infinite space of possible vocabularies. We anticipate these results may further our understanding of problems in AI safety and general intelligence.

### v4

@@ 3 Limits of Intelligence
As stated earlier, intelligence is often understood in terms of “the ability to generalise” [chollet2019] and learn the policy which “caused” examples [bennett2023c]. We wish to understand the objective upper limits of intelligence. $<_{w}$ is the optimal choice of proxy for sample efficiency [bennett2023b, prop. 1, 2, 3]. In the context of a fixed abstraction layer, the upper bound of intelligent behaviour is attained by the weakness proxy. However, more intelligence is not always useful. Intuitively, making a human more intelligent is unlikely to improve their driving. Likewise, intelligence conveys no advantage if one’s purpose is to ascribe meaning to random noise. The extent to which it is possible to construct weak correct policies depends on both the abstraction layer and the task. An abstraction layer is a bottleneck on intelligence, and can be goal directed just like a task. We can measure this goal directed “utility” as the extent to which it is possible to construct weaker correct policies given a task in a particular abstraction layer.
@@
Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
To maximise probability of generalisation, we use the weakness proxy and try to maximise utility. Increasing utility changes the abstraction layer to allow construction of weaker correct policies. Beyond a certain point, utility may be increased by increasing $\lvert L_{\mathfrak{v}}\rvert$ without actually changing the weakest correct policies $\Pi_{\gamma}$ contains (just the size of their extensions). This means it is not always necessary to increase utility to construct policies that generalise, but it is helpful up to a point. Utility is maximised when $\mathfrak{v}=P$, though in practice finite resources would limit us to smaller vocabularies. $\Gamma_{P}$ contains all tasks in all vocabularies. Hence, for every task $\rho$ in $\Gamma_{P}$ we can define a function that takes a vocabulary $\mathfrak{v}$ and returns a $\mathfrak{v}$-task which is a child of $\rho$. This lets us represent a task in different vocabularies to compare their utility.
Definition 7 (uninstantiated-tasks)
The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a $\mathfrak{v}^{\prime}$-task $\omega\sqsubset\rho$. We call $\lambda_{\rho}$ an uninstantiated-task. It is instantiated by choosing a vocabulary.
Proposition 1 (upper bound)
The most ‘intelligent’ choice of policy and vocabulary given uninstantiated task $\lambda_{\rho}$ is $\pi$ and $\mathfrak{v}$ s.t. $\mathfrak{v}$ maximises utility for $\lambda_{\rho}(\mathfrak{v})$, $\pi\in\Pi_{\lambda_{\rho}(\mathfrak{v})}$ and $\pi$ maximises weakness.
Proof
We have equated intelligence with sample efficient generalisation. According to [bennett2023b, prop. 1,2] the weakest correct policies have the highest probability of generalising. Given an uninstantiated task $\lambda_{\rho}$, utility to measures the weakness of the weakest correct policies in different vocabularies. By choosing $\mathfrak{v}$ which maximises utility for $\lambda_{\rho}(\mathfrak{v})$, we instantiate $\lambda_{\rho}$ in a vocabulary that maximises the weakness of correct policies for $\lambda_{\rho}$. Then, using weakness proxy, we can select a policy that has the highest possible probability of generalising, and thus maximise sample efficiency. ∎
Put another way, utility is maximised for $\lambda_{\rho}(\mathfrak{v})$ when $\mathfrak{v}=P$. It follows that when utility is maximised there must exist $\pi\in\Pi_{\lambda_{\rho}(\mathfrak{v})}$ which is a weakest policy in $\Pi_{\lambda_{\rho}(P)}$. This is why maximising utility maximises the weakness of correct policies.
Concluding remarks: Here we proposed computational dualism, and presented an argument in support of an alternative which allows for objective claims regarding behaviour. We proposed an objective upper bound for intelligent behaviour. In general this upper bound is unattainable, because we can only build systems with finite vocabularies. Moreover, too large a vocabulary $\mathfrak{v}$ could make $\mathfrak{v}$-tasks intractable. However in practical terms higher utility does mean weaker and more generalise-able policies, which suggests one should optimise for higher utility whilst trying to minimise the size of the vocabulary. We anticipate these results may further our understanding of AI safety and general intelligence.
[…]